OpenStudy (anonymous):

To solve the equation x^2 +6x=1 by the technique of completing the square, you would add 9 to each side of the equation as one of the steps to reaching the solution. True False

OpenStudy (anonymous):

If you add 9 to both sides, you would get \[x^2 + 6x + 9 = 10\] Can this; \[x^2 + 6x + 9 = 10\] be converted into a squared form(the form of two binomials)? Now, take the square root of constant number ( 9), add it to x, and square that. You would then get \[(x + 3)^2\]. So the answer would be true. To check, we can use FOIL.\[(x + 3)(x + 3) = x^2 + 3x + 3x + 9 = x^2 + 6x + 9\]

OpenStudy (anonymous):

OMG! thank you so much! Your the 2nd person who's helped me! I really appreciate it.

OpenStudy (anonymous):

yw :)